1.Assume that an appliance manufacturer is employing variable factors X and Y in such amounts that the MRPs of the last units of X and Y employed are $100 and $60 respectively. Factor X can be hired at $50 per unit and factor Y at $20 per unit. If you are the manager of the firm, what is your proper employment strategy for maximizing your firm’s profit?

2.Assume that labour and capital are substitutes in production. If there is an increase in the price of capital, how can this lead to either an increase or decrease in the demand for labour?
Answer #1

1) at profit maximisation, MRTS= MRP1/MRP2 = P1/P2

MRP1/MRP2= 100/60= 1.67

P1/P2= 50/20= 2.5

Since MRP1/MRP2 is less than P1/P2, firm should less of X and more of Y to get both the ratios equal, in order to maximise profit.

2) if labor and capital are substitutes and if price of capital increases such that marginal rate of substitution is greater than price ratio of labor and capital, increase in price of capital leads to increase in demand for labor.